OpenStudy (dumbsearch2):

Eight hours after a walker and jogger leave the house at the same time and head in the same direction, the jogger is 136 miles ahead of the walker. The jogger's speed is 10 mph more than 2 times the speed of walker's. Find the speed of the jogger.

OpenStudy (mathstudent55):

The walker walks at w mph The jogger jogs at j mph We know that j = 2w + 10 d = rt, distance = rate (speed) x time The distance the walker covers in 8 hours is: d = w*8 = 8w The distance the jogger covers in 8 hours is d = (2w + 10)*8 = 16w + 80 The jogger covered 136 more miles in the same time, so 16w + 80 = 8w + 136 Solve for w to get the speed of the walker. Then substitute the value of w into 2w + 10 to get the speed of the jogger.